Circular No. 5755

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRB 930309 T. E. Harrison and B. McNamara, Astronomy Department, New Mex-

ico State University, on behalf of the Comptel/NMSU Rapid Response Network, report: “The updated coordinates for the variable radio source discovered within the Comptel error box for GRB 930309 (IAUC 5750) are nearly coincident with the position of the IRAS source 21311+5426, whose 12-, 25-, and 60-μm fluxes are consistent with a 400-K blackbody spectral energy distribution. There are no optical counterparts visible on the POSS at this position, though a number of faint stars are located within 1′ of it. We encourage observations at other wavelengths, especially near-infrared images, to attempt to determine the nature of this variable radio source. Interested observers can contact the Comptel/NMSU Rapid Response Network (e-mail bmcnamar@nmsu.edu or jryan@comptel.unh.edu).”

SUPERNOVA 1993J IN NGC 3031 D. King reports: “Observations of the Na I D_{2},D_{1} doublet were ob-

tained on the night of Apr. 5/6 with the Utrecht échelle spectrograph (resolving power 48 500) on the William Herschel Telescope at La Palma. Six components were found as follows: –139 km/s (weak), –123 km/s (strong), –57 km/s (weak), –3 km/s (strong), +119 km/s (strong), +133 km/s (strong).”

M. Richmond writes that he and A. Filippenko, B. Leibundgut, Y. Paik, and R. R. Treffers (University of California at Berkeley) confirm that SN 1993J has brightened. CCD images taken with the Leuschner 0.76-m telescope reveal the following differences between the supernova’s magnitude on Apr. 6.2 and 7.2 UT: U, +0.05; B, –0.07; V, –0.09; R, –0.13; I, –0.06. Spectral measurements are needed to help interpret this brightening.

Further CCD photometry (cf. IAUC 5750) reported by T. Kato (K) at Ouda Station, Nara, and by M. Okyudo and T. Ishida (O) at Nishi-Harima Observatory, all obtained using Corwin’s values for GSC 4383.0928 on IAUC 5742: Apr. 6.485 UT, V = 11.84 (K); 6.538, B = 12.23 (K); 6.511, I_{c} = 11.29 (K); 6.685, V = 11.79 (K); 6.720, B = 12.19 (K); 6.739, I_{c} = 11.23 (K); 7.496, V = 11.70 (K); 7.532, B = 12.11 (K); 7.567, I_{c} = 11.18 (K); 7.589, V = 11.73 (K); 7.610, V = 11.69 (O).

1993 April 8 Daniel W. E. Green